kekeman:

Ellis is painting wooden fenceposts before putting them in his yard. They are each 7 feet tall and have a diameter of 1 foot. There are 11 fenceposts in all. How much paint will Ellis need to paint all the surfaces of the 11 fenceposts? Use 3.14 for π, and round your answer to the nearest hundredth. Provide an explanation and proof for your answer to receive full credit.

kekeman:

Help plss

new2luv:

is there an image

kekeman:

Nope

new2luv:

give me a second @AZ

AriahMHebrank:

we need to know how much surface area the paint will cover but here is how to find that surface area Step-by-step explanation: 1 fence post has a surface area of pi*diameter* height pi * 1 * 5 =15.707963 10* fence post =157.07963 sq ft

kekeman:

Are u sure

AZ:

That's wrong

AZ:

I remember answering this exact question a few weeks ago too Surface area of cylinder = \( 2\pi rh+2\pi r^2 \) So plug in r = 0.5 diameter is 1 so the radius is half of that and height is 7 so what is the surface area of 1 wooden fencepost?

kekeman:

259.05 sq

AZ:

That would be your final answer, yes you multiply the surface area by 11 so you can find how much surface area you're painting for 11 fenceposts However, keep in mind that the units should be square feet!

kekeman:

Yess ok